# Optimizing Crowd-Aware Multi-Agent Path Finding through Local Broadcasting with Graph Neural Networks

### v2

@@ I INTRODUCTION
On the contrary, decentralized methods [15, 16, 17, 18, 19, 1] seek to tackle these challenges by allowing each agent to acquire their own policies. In these approaches, agents can reactively plan paths within partially observable environments. Such methods prove beneficial in situations where agents lack comprehensive knowledge of the world, as is often the case in the domain of autonomous vehicles. Rather than pursuing an optimal solution for all agents, decentralized planners train local policies that rapidly generate sub-optimal solutions, as a tradeoff between speed and solution quality. Given that agents make their decisions based on local information, decentralized approaches often face challenges in achieving effective global coordination among agents. In cluttered, dynamic environments characterized by congestion or rapid changes, agents may tend to prioritize their individual goals, potentially resulting in conflicts and inefficiencies that affect the overall system’s performance.
@@
To tackle the challenges presented by dense crowds, our approach, CRAMP, integrates a crowd-awareness mechanism and local communication via Graph Neural Networks (GNN). This strategy is guided by a boosted curriculum training approach, which gradually exposes agents to increasingly complex scenarios. CRAMP is specifically designed to train intelligent agents capable of efficiently navigating in dense and dynamic environments. By formulating the MAPF problem as a sequential decision-making task, we employ deep reinforcement learning techniques, enabling agents to make optimal decisions while considering the presence of other agents and the dynamic nature of the environment.
@@ II RELATED WORK
Moreover, the exploration of hybrid techniques that integrate both centralized and decentralized components remains an important avenue of research [23, 24]. Hybrid approaches could potentially capitalize on the strengths of both paradigms while mitigating their respective limitations. In this work, we primarily study the centralized and decentralized approaches. Future research could explore hybrid approaches to gain a more comprehensive understanding of the trade-offs involved and to further advance the field.
@@ II-D Graph neural networks for MAPF
Despite the promising potential, limited work exists in applying GNNs to MAPF. Notably, [40] integrated GNNs into MAPF but trained the model through imitation learning, leading to an imbalance between exploration and exploitation. As a consequence, the model’s generalization in new environments is compromised. In this work, we aim to address this issue by integrating GNNs into our framework, seeking to enhance the generalization capabilities of our model. Our contribution to this landscape includes a novel crowd-avoidance mechanism, supported by local communication through GNNs. We adopt a simple yet effective curriculum-driven training strategy to enhance learning efficiency and significantly reduce incidents of collisions and deadlocks.
@@ III-B Reinforcement learning for local policy training
We train the agent’s policy using deep reinforcement learning and demonstration learning techniques. Multiple agents are independently trained in a shared environment utilizing the Asynchronous Advantage Actor-Critic (A3C) algorithm [45]. The inputs to the policy network comprise four layers of an agent’s local observation, which include the grid world, obstacles, neighboring agents, and neighbors’ goals, alongside its own goal position. The network’s outputs consist of the next action, predicted return value, validity of the action, and blocking probability. The detailed architecture of the policy network is illustrated in Fig. 2. The CNN feature extractor employs two sequential blocks of [Conv2D-Conv2D-Conv2D-Relu-MaxPool]. The CNN’s output is a $512\times 1$ vector. This feature vector is subsequently input into an LSTM to encode temporal features. For the GNN block, the adjacency matrix is computed from the local field of view (FOV). For each agent within the ego agent’s FOV, encoded feature vectors are extracted to create a node embedding matrix of size $N\times 512$, where $N$ represents the maximum number of visible agents. The GNN processes both the adjacency and node embedding matrices to produce an $N\times 128$ matrix, with each row reflecting a graph signal from each agent. These nodes are averaged and concatenated with the LSTM feature vector. The resulting concatenated feature vector is fed into fully-connected layers to predict the variables.
@@
The policy is updated by adding the entropy term $H=-\sum_{a}\pi(a_{t}=a|o_{t})\log(\pi(a_{t}=a|o_{t}))$ to the policy loss:
@@ III-D Crowd-aware reward function
Decentralized policies might result in agents behaving selfishly, blocking others from their goals, and potentially causing deadlocks in tight areas. To foster teamwork, agents receive a significant reward when all complete an episode together, but face penalties for hindering others, promoting cooperation and efficiency in the multi-agent system.
@@
The threshold $\zeta$ is not fixed; its magnitude should be higher when the environment contains more agents and obstacles and lower in sparser environments. We define $\zeta$ as follow:
@@ IV-C Result analysis
As depicted in Table I, our model consistently demonstrates better performance across most environments, particularly in terms of success rate. Notably, our approach excels in resolving challenges within two densely populated scenarios, featuring 32 agents with a 0.3 density and 64 agents with a 0.2 density, where all other methods encounter complete failure. Across the remaining configurations, our model consistently either secures the top position or closely follows the best-performing method.

### v3

# Optimizing Crowd-Aware Multi-Agent Path Finding through Local Communication with Graph Neural Networks

@@ I INTRODUCTION
On the contrary, decentralized methods [15, 16, 17, 18, 19, 1] seek to tackle these challenges by allowing each agent to acquire their own policies. In these approaches, agents can reactively plan paths within partially observable environments. Such methods prove beneficial in situations where agents lack comprehensive knowledge of the world, as is often the case in the domain of autonomous vehicles. Rather than pursuing an optimal solution for all agents, decentralized planners train local policies that rapidly generate sub-optimal solutions as a tradeoff between speed and solution quality. Given that agents make their decisions based on local information, decentralized approaches often face challenges in achieving effective global coordination among agents. In cluttered, dynamic environments characterized by congestion or rapid changes, agents may tend to prioritize their individual goals, potentially resulting in conflicts and inefficiencies that affect the overall system’s performance.
@@
To tackle the challenges presented by dense crowds, our approach, CRAMP, integrates a crowd-awareness mechanism and local communication via Graph Neural Networks (GNN) [20]. This strategy is guided by a boosted curriculum training approach, which gradually exposes agents to increasingly complex scenarios. CRAMP is specifically designed to train intelligent agents capable of efficiently navigating in dense and dynamic environments. By formulating the MAPF problem as a sequential decision-making task, we employ deep reinforcement learning techniques, enabling agents to make optimal decisions while considering the presence of other agents and the dynamic nature of the environment.
@@ II RELATED WORK
Moreover, the exploration of hybrid techniques that integrate both centralized and decentralized components remains an important avenue of research [24, 25]. Hybrid approaches could potentially capitalize on the strengths of both paradigms while mitigating their respective limitations. In this work, we primarily study the centralized and decentralized approaches. Future research could explore hybrid approaches to gain a more comprehensive understanding of the tradeoffs involved and to further advance the field.
@@ II-D Graph neural networks for MAPF
Despite the promising potential, limited work exists in applying GNNs to MAPF. Notably, [41] integrated GNNs into MAPF but trained the model through imitation learning, leading to an imbalance between exploration and exploitation. As a consequence, the model’s generalization in new environments is compromised. In this work, we aim to address this issue by integrating GNNs into our framework, seeking to enhance the generalization capabilities of our model. Our contribution to this landscape includes a novel crowd-avoidance mechanism supported by local communication through GNNs. We adopt a simple yet effective curriculum-driven training strategy to enhance learning efficiency and significantly reduce incidents of collisions and deadlocks.
@@ III-B Reinforcement learning for local policy training
We train the agent’s policy using deep reinforcement learning and demonstration learning techniques. Multiple agents are independently trained in a shared environment utilizing the Asynchronous Advantage Actor-Critic (A3C) algorithm [46]. The inputs to the policy network comprise four layers of an agent’s local observation, which include the grid world, obstacles, neighboring agents, and neighbors’ goals, alongside its own goal position. The network’s outputs consist of the next action, predicted return value, validity of the action, and blocking probability. The detailed architecture of the policy network is illustrated in Fig. 2. The CNN feature extractor employs two sequential blocks of [Conv2D-Conv2D-Conv2D-Relu-MaxPool]. The CNN’s output is a $512\times 1$ vector. This feature vector is subsequently input into an LSTM to encode temporal features. In the GNN block, the adjacency matrix is computed based on the local field of view (FOV), where the ego agent is connected to all other agents within the FOV. For each agent within the ego agent’s FOV, encoded feature vectors are extracted to create a node embedding matrix of size $N\times 512$, where $N$ represents the maximum number of visible agents. The GNN processes both the adjacency and node embedding matrices to produce an $N\times 128$ matrix, with each row reflecting a graph signal from each agent. These nodes are averaged and concatenated with the LSTM feature vector. The resulting concatenated feature vector is fed into fully-connected layers to predict the variables.
@@
The policy is updated by adding the entropy term $H=-\sum_{a}\pi(a_{t}=a|o_{t})\log(\pi(a_{t}=a|o_{t}))$ [47] to the policy loss:
@@ III-D Crowd-aware reward function
Decentralized policies might result in agents behaving selfishly, blocking others from their goals, and potentially causing deadlocks in tight areas. To foster teamwork, agents receive a significant reward when all complete an episode together but face penalties for hindering others, promoting cooperation and efficiency in the multi-agent system.
@@
The threshold $\zeta$ is not fixed; its magnitude should be higher when the environment contains more agents and obstacles and lower in sparser environments. We define $\zeta$ as follows:
@@ IV-C Result analysis
As depicted in Table I, our model consistently demonstrates better performance across most environments, particularly in terms of success rate. Notably, our approach can find solutions in two densely populated scenarios, featuring 32 agents with a 0.3 density and 64 agents with a 0.2 density, where all other methods encounter complete failure. Across the remaining configurations, our model consistently either secures the top position or closely follows the best-performing method.